Click Here to receive Service Notes automatically via email.
Date | Description |
---|---|
07/13/2007 | Network Drives (W:, Y: and U:) were briefly unavailable this morning due to an electrical power outage in the McClain Hall data center.
Electrical power has been restored and network drives are back up. |
07/03/2007 | The phone lines feeding the Rec Center have been repaired. Phone service should be back to normal in this building.
Telephone services will continue to work today to repair damaged lines feeding Centennial Hall. |
07/03/2007 | Contractors performing excavation work South of the Student Union building have damaged a large telephone feeder cable. Telephone service to Centennial Hall and the Student Rec Center has been impacted. Telephone Services is working to restore service as soon as possible. |
06/15/2007 | Due to the renovation work taking place in Brewer Nason and Blanton halls it was requested we shutdown all network equipment. Network services have been shutdown in Brewer Nason and Blanton as requested. |
06/15/2007 | Due to Fair Apartment being unoccupied and high temperatures in the Fair network equipment closets, network services to Fair Apartments have been shutdown for the summer. |
06/08/2007 | The cable TV Weather channel will be unavailable until Friday due to roofing repairs on the Student Union Building. |
06/08/2007 | The cable TV CNN channel will be unavailable until Friday due to roofing repairs on the Student Union Building. |
06/04/2007 | The phone lines between Violette Hall and McKinny Center have been cut due to work being conducted on a sewer line. Repairs are being made as rapidly as possible. As of approximately 10:30 the operator service was restored. Voice mail is still not functional. |
06/04/2007 | Phone services to the operator and voice mail have been disrupted. The cause for this disruption of serivce is being investigated. |
06/01/2007 | Banner archive logs ran out of filesystem space and hung all Banner ADMIN services until space was made available. The system hung at 5:27am and was available again at 8:22am. Work with the new ODS integration with Banner is believe to be causing the significant increase in archive logs. Steps have been put into place to deal with the increased logging and to monitor it more closely. |
05/18/2007 | Date Correction (5/18 not 5/17): TruView will be restarted at about 6:00 a.m. on Friday 5/18/07. This is during the monthly scheduled maintenance window (which runs from 8:00 p.m. - 6:00 a.m. on the third Thursday) and during a period of low usage. |
05/15/2007 | MOREnet has corrected the problem in the state network. Our Internet access is back. |
04/30/2007 | The brief Internet outage earlier today was caused by a MOREnet state network outage in Kansas City. Internet services have been restored. |
04/27/2007 | TruView was unavailable for approximately 15-20 minutes starting at 2:00 p.m. on 4/27/2007. The error was caused by a server issue which has been resolved. All services have been restored. |
04/23/2007 | The Truman network problems have been resolved. Multiple network services were not performing normally. Those problems have been resolved. All network services are back up. We are reviewing the cause of the problems. |
04/19/2007 | Tomorrow evening (Thur. 4/19) during the scheduled maintenance window, an update will be applied to the PureMessage spam and virus filter portion of the email server. The actual server down time is estimated to be less than 1 hour. |
04/19/2007 | Research in Motion (RIM) has restored BlackBerry service. |
04/16/2007 | The organization web server (neon) was rebooted at approximately 4:00 p.m. due to slow response times on dynamic websites. The reboot has fixed the problem and all sites appear to be back online. |
04/16/2007 | All Banner services were down this morning due to a full filesystem where the system's archive logs are written. The filesystem space issue has been addressed by freeing up space in the filesystem and appropriate modifications have been made to the automated script that manages this. |
04/06/2007 | The Exchange email system is back up. Full faculty / staff email services are back up. |
04/04/2007 | Campus cable TV services were briefly unavailable this afternoon due to an electrical power outage in the Student Union Building. Cable TV service has been restored. |
04/01/2007 | SBT Maintenance to address security was successful. |
03/31/2007 | A vulnerability in the SBT software has been identified. In the interest of security we have locked down editing on the website for approximately 48 hours to deploy and apply a patch that will eliminate the vulnerability. |
03/23/2007 | A core network switch is failing. Twenty six servers are connected to the failing switch. All IT services are still up but the failing switch component needs replaced before it does affect IT services.
We've received the necessary replacement parts and will install them during the maintenance window. We are planning to do a phased replacement and switch the servers over one at a time to the new component. Each server should be down only 1-2 seconds during the switchover. |
03/21/2007 | Due to a configuration change in Banner, the opening of today's fall registration group at 5pm resulted in students getting an error stating it wasn't their time to register. The configuration change was backed off and enrollment ability immediately opened up. Normal operation was back by approximately 5:07pm. |